### v2

# Doracamom: Joint 3D Detection and Occupancy Prediction with Multi-view 4D Radars and Cameras for Omnidirectional Perception

Abstract
3D object detection and occupancy prediction are critical tasks in autonomous driving, attracting significant attention. Despite the potential of recent vision-based methods, they encounter challenges under adverse conditions. Thus, integrating cameras with next-generation 4D imaging radar to achieve unified multi-task perception is highly significant, though research in this domain remains limited. In this paper, we propose Doracamom, the first framework that fuses multi-view cameras and 4D radar for joint 3D object detection and semantic occupancy prediction, enabling comprehensive environmental perception. Specifically, we introduce a novel Coarse Voxel Queries Generator that integrates geometric priors from 4D radar with semantic features from images to initialize voxel queries, establishing a robust foundation for subsequent Transformer-based refinement. To leverage temporal information, we design a Dual-Branch Temporal Encoder that processes multi-modal temporal features in parallel across BEV and voxel spaces, enabling comprehensive spatio-temporal representation learning. Furthermore, we propose a Cross-Modal BEV-Voxel Fusion module that adaptively fuses complementary features through attention mechanisms while employing auxiliary tasks to enhance feature quality. Extensive experiments on the OmniHD-Scenes, View-of-Delft (VoD), and TJ4DRadSet datasets demonstrate that Doracamom achieves state-of-the-art performance in both tasks, establishing new benchmarks for multi-modal 3D perception. Code and models will be publicly available.
Index Terms:
Autonomous driving, camera, 4D radar, deep learning, omnidirectional perception, 3D object detection, 3D occupancy prediction.
I Introduction
Figure 1: Performance comparison of different methods on 3D perception benchmarks. By effectively fusing camera and 4D radar inputs, Doracamom (marked with $\star$) consistently achieves superior performance across multiple metrics, outperforming existing camera-only and camera-radar fusion approaches in both 3D object detection and occupancy prediction tasks.
Autonomous driving technology is at the forefront of modern transportation revolution, attracting considerable attention. Autonomous driving systems typically include components such as environmental perception, trajectory prediction, and planning control to achieve self-driving capabilities. Accurate 3D perception is a critical foundation, focusing mainly on 3D object detection and semantic occupancy prediction tasks. 3D object detection employs 3D bounding boxes to locate foreground objects in the scene and predicts attributes such as category and velocity, and belongs to sparse scene representation [1]. In contrast, semantic occupancy uses fine-grained voxel representation to capture the geometric and semantic features of a scene, which is a form of dense scene representation [2]. To accomplish these tasks, sensors such as onboard cameras, LiDAR, and millimeter-wave radars are commonly used to collect environmental data as input.
Among these sensors, LiDAR operates on the Time of Flight (TOF) principle, emitting and receiving laser beams to generate dense point clouds, providing high-precision geometric representations of the environment [3, 4, 5, 6, 7]. However, LiDAR is susceptible to adverse weather and is costly [8]. In contrast, cameras and radars are more cost-effective, making them suitable for large-scale deployment. Cameras capture rich color and texture information with high resolution but lack depth information and are vulnerable to weather interference [9]. Radar, on the other hand, emits electromagnetic waves to detect target distance, Doppler, and scattering information, offering robustness against weather conditions [10].
The 4D imaging radar, an advancement over conventional radar, not only includes additional elevation information but also provides higher resolution point cloud than conventional $2+1$D radar. Recent studies have shown that it has considerable promise in a various downstream tasks [11, 12]. However, compared to LiDAR, its point cloud remains sparse and noisy. Thus, cross-modal fusion is essential to effectively compensate for these shortcomings, emphasizing the need to integrate information from cameras and 4D radar.
In recent years, especially with the advent of 4D radar datasets, research on 4D radar and camera fusion has demonstrated significant potential in the field of perception. Most mainstream fusion techniques now employ a Bird’s Eye View (BEV) architecture, transforming raw sensor inputs into BEV features for integration. This perspective helps to mitigate occlusion issues with foreground objects and maintains scale consistency. The essence of these methods is the conversion of image data, which lacks depth information, into a bird’s eye view using advanced view transformation algorithms. For occupancy prediction task, most research focuses on vision-centered or vision and LiDAR fusion, as occupancy prediction requires fine-grained voxel representation and semantic information. Traditional radar, lacking height information, is not suitable for 3D occupancy prediction. In contrast, 4D radar offers new possibilities by accessing elevation information and higher resolution point cloud, though related research is still limited. In addition, integrating 3D object detection and occupancy prediction as two key perception tasks within a unified multi-task framework can optimize computational resources and efficiency, providing substantial benefits.
To this end, we introduce Doracamom, a unified framework that, for the first time, fuses multi-view cameras with 4D radar point clouds to handle 3D object detection and semantic occupancy prediction tasks simultaneously. The main contributions of this paper are concluded as below:
  • We propose Doracamom, which is the first unified framework to fuse cameras and 4D radar for joint 3D object detection and occupancy prediction, enabling comprehensive environmental perception and understanding.
  • Specifically, we design three key components to enhance model performance. The coarse voxel query generator (CVQG) leverages 4D radar geometric cues and image semantic information to establish well-initialized voxel queries for effective feature refinement. Dual-branch temporal encoder (DTE) enables parallel temporal modeling in both BEV and voxel domains, capturing comprehensive spatio-temporal representations of the scene. A cross-modal BEV-voxel fusion (CMF) module adaptively fuses complementary features through attention mechanisms, incorporating auxiliary binary occupancy prediction and BEV segmentation tasks to guide the feature learning process for more discriminative representations.
  • The comprehensive experiments show that Doracamom achieves state-of-the-art performance and makes new benchmark for both 3D object detection and occupancy prediction with 4D radar and camera fusion, on several 4D radar datasets including OmniHD-Scenes [13], VoD [14], and TJ4DRadSet [15], as shown in Fig. 1.
II Related Works
II-A 3D Perception with Camera
Recent 3D perception research has primarily focused on vision-based approaches. While early work directly regressed 3D attributes from single images [16], such monocular methods suffer from occlusion and viewpoint changes. This has led to increased interest in multi-view approaches that transform 2D features into 3D space [17, 18, 19, 20, 21] using BEV representations.
Existing methods can be categorized into three main approaches. The first uses depth prediction, starting with LSS [18] which lifts 2D features to BEV space through depth distributions and voxel pooling. This was extended by BEVDepth [20], BEVDet [19], and BEVDet4D [22] with LiDAR supervision and temporal fusion [23]. The second approach employs back projection, where methods like OFTNet [24] and simpleBEV [25] project 3D voxels onto 2D images for feature sampling. The third approach is attention-based, where BEVFormer [21] uses BEV queries and deformable attention [26] for adaptive feature aggregation, while DETR3D [27], PETR [28], and Sparse4D [29] demonstrate strong performance using object queries without explicit BEV features.
Vision-based occupancy prediction transforms 2D features to 3D space to obtain BEV features [30], TPV features [31], or voxel features [32]. Methods are categorized into projection-based, depth-based, and cross-attention-based approaches [2]. MonoScene [33] uses projection and U-Net for semantic completion, while FlashOcc [30] and FastOcc [34] follow LSS [18] to predict depth distributions, using channel-to-height operations [30] for memory efficiency. SurroundOcc [32], PanoOcc [35], and TPVFormer [31] employ deformable attention [26] for feature aggregation. Recent works include AdaptiveOcc [36], using octree for voxel representation, and LinkOcc [37], incorporating sparse queries and temporal association through near-online training and contrastive learning.
II-B 3D Perception with Traditional Radar and Camera Fusion
Conventional automotive radar is constrained by its angular resolution, resulting in sparse point clouds lacking height information, necessitating camera-radar fusion for enhanced perception [10]. CenterFusion [38] generates frustum ROIs using 2D detectors and associates them with radar pillars. Feature-level fusion methods work at either BEV level [39, 40] or proposal stage [41, 42]. CRN [39] uses radar for BEV transformation and multi-modal deformable attention for feature alignment. RCBEV [43] combines point fusion and ROI fusion, while RCBEVDet [40] employs RadarBEVNet with dual representations and RCS- aware BEV features, employing a cross-modal attention mech- anism to merge radar and image BEV features. CRAFT [41] associates proposals with radar points in polar coordinates, while CramNet [44] uses ray-constrained attention for geometric correspondence. TransCAR [42] and FUTR3D [45], both extending DETR3D [27], utilize object queries for radar-camera feature interaction.
Beyond 3D detection, recent works explore camera-radar fusion for occupancy prediction. Occfusion [46] employs dynamic 2D/3D fusion for multi-level feature representation, while LiCROcc [47] uses cross-modal distillation modules for semantic scene completion. However, the sparsity and lack of height information in traditional radar remain challenging for 3D semantic occupancy tasks.
II-C Recent Progress on Perception with 4D Radar
The 4D imaging radar offers significant improvements over conventional radar in resolution and point cloud density, along with elevation resolution, greatly expanding its potential applications [11, 12]. The recent emergence of 4D radar datasets [48, 15, 14, 49, 50, 51, 13] has spurred research in this field. Some studies focus on leveraging the characteristics of 4D radar for multi-object tracking [52, 53]. Other researches utilize 4D radar point clouds for mapping and localization [54, 55, 56], scene flow estimation [57, 58], panoptic segmentation [59, 60] and visual grounding [61, 62, 63], among other applications.
In the field of 3D object detection, many research efforts have achieved remarkable results by using 4D radar or fusing other sensors. RPFA-Net [64] uses self-attention for 4D radar feature extraction, while SMURF [65] mitigates the sparsity and noise of 4D radar point clouds using pillarization and density features derived from Kernel Density Estimation (KDE) of multi-dimensional Gaussian mixtures. SCKD [66] employs semi-supervised knowledge distillation from LiDAR-radar fusion teacher networks. InterFusion [67] employs a self-attention mechanism to learn features from both LiDAR and 4D radar modalities, exchanging information at intermediate layers. RCFusion [68] leverages RadarPillarNet for hierarchical feature extraction from 4D radar point clouds and effectively fuses BEV features of camera and 4D radar modalities using interactive attention. UniBEVFusion [69] introduces the Radar Depth LSS (RDL) module to improve depth estimation and employs Unified Feature Fusion (UFF) to integrate features from different modalities. LXL [70] proposes a ”radar occupancy-assisted depth-based sampling” strategy that assists image view transformation by combining predicted depth with 3D radar occupancy grids. SGDet3D [71] proposes dual-branch fusion with geometric and semantic information processing, coupled with object-oriented attention for effective radar-camera feature interaction, while HGSFusion [72] introduces hybrid point generation and dual synchronization for radar-camera fusion. Recently, DPFT [73] fuses the 4D radar tensors with camera feature by sampling from queries in 3D space, achieves competitive performance.
For occupancy prediction tasks, RadarOcc [74] achieves binary occupancy prediction using 4D radar tensors, in which only foreground and background are distinguished without providing more specific semantic information. Considering the significant resources consumption required for processing raw radar tensors, such approach is impractical for real-time applications. On the other hand, radar point clouds are widely accepted in ADAS and autonomous driving systems, yet research on 3D semantic occupancy prediction using practical 4D radar point clouds remains limited.
III Methodology
III-A Overall Architecture
We present Doracamom, a unified multi-task framework that fuses multi-view images and 4D radar point clouds for joint 3D object detection and occupancy prediction. The overall architecture is shown in Fig. 2.
Figure 2: The overall framework of Doracamom. Initially, the camera encoder and 4D radar encoder are utilized to extract multi-camera features and radar BEV features, respectively. Subsequently, the coarse voxel query generator employs geometric priors derived from 4D radar features and semantic priors from image features to generate coarse voxel queries. These queries are then refined by a stacked voxel queries encoder to obtain local fine-grained voxel features. A dual-branch temporal encoder is employed to fuse historical BEV and voxel features with current frame features, leveraging temporal clues. The output radar BEV and image voxel features are fed into the cross-modal BEV-voxel fusion module for adaptive fusion, resulting in the final BEV and voxel representations. Finally, the obtained representations are used to predict 3D detection and semantic occupancy for the current scene.
Initially, multi-view images and 4D radar point clouds are fed into the camera and 4D radar encoders to extract image 2D features and 4D radar BEV features, respectively. These features are then passed to the coarse voxel query generator, which combines image and radar features to generate geometrically-semantic-aware coarse-grained voxel queries. The voxel query encoder iteratively enhances fine-grained voxel features through stacked transformer blocks using cross-view attention. Subsequently, a dual-branch temporal encoder leverages temporal cues to enhance both BEV and voxel feature representations. The cross-modal BEV-voxel fusion module adaptively integrates the feature representations from both modalities, obtaining the final BEV and voxel features, which are then fed into the multi-task head for prediction.
The details of the proposed method will be elaborated in the following subsections.
III-B Camera & 4D Radar Encoders
In the feature extraction stage, we employ a decoupled architecture for independent high-dimensional feature extraction from two input modalities. The camera encoder processes multi-view images represented as $I\in\mathbb{R}^{N_{C}\times 3\times H_{I}\times W_{I}}$, where $N_{C}$ is the number of cameras, $H_{I}$ and $W_{I}$ are the image height and width, and 3 corresponds to the RGB channels. Feature extraction is performed using a shared ResNet-50 [75] backbone network and a Feature Pyramid Network (FPN) [76] as the neck structure, which obtains multi-scale features and simplifies them into a single-scale representation $\mathcal{F}_{I}=\{F_{I}^{i}\}_{i=1}^{N_{C}}$, where $F_{I}^{i}\in\mathbb{R}^{C\times H_{C}\times W_{C}}$ is the feature of the $i$-th view, $C$ denotes the feature dimension, and $H_{C}$ and $W_{C}$ are the spatial dimensions of the feature map.
To address the sparsity issue of 4D radar point clouds and obtain their ground velocities by eliminating ego-motion effects, we implement a pre-processing pipeline that combines multi-frame radar point clouds accumulation and velocity compensation. The algorithm processes each radar sweep $s$ using the corresponding ego vehicle velocity $\mathbf{V}_{e}\in\mathbb{R}^{3\times 1}$, which is transformed into the radar coordinate system through the radar-to-ego rotation matrix $\mathbf{R}_{r\to e}\in\mathbb{R}^{3\times 3}$, yielding $\mathbf{V}_{rad}\in\mathbb{R}^{3\times 1}$. To compensate for relative radial velocity, the velocity vector is decomposed into the radial direction based on each point’s azimuth angle $\phi$ and elevation angle $\theta$. The compensated velocity components are then transformed to the current ego coordinate system using rotation matrices $\mathbf{R}_{r\to e}\in\mathbb{R}^{3\times 3}$ and $\mathbf{R}_{e\to e^{\prime}}\in\mathbb{R}^{3\times 3}$. For each point’s position, the transformation is achieved using the radar-to-ego transformation matrices $\mathbf{R}_{r\to e}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}_{r\to e}\in\mathbb{R}^{3\times 1}$, along with the ego pose transformation matrices from sweep time to current time $\mathbf{R}_{e\to e^{\prime}}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}_{e\to e^{\prime}}\in\mathbb{R}^{3\times 1}$. Beware the motion of points which introduced by motion of surrounding dynamic objects are omitted during the accumulation operation, since such motions rarely introduce large errors.
The 4D radar encoder processes input point clouds $\mathcal{P}\in\mathbb{R}^{N_{R}\times D}$, where $N_{R}$ and $D$ represent the number of points and attribute dimension, respectively. We adopt RadarPillarNet [68] to encode the input 4D radar point clouds, which generates pseudo images through hierarchical feature extraction. The encoded features are then processed by SECOND and SECONDFPN [77] to produce refined 4D radar BEV features $\mathcal{F}_{R}\in\mathbb{R}^{C_{R}\times H\times W}$, with $C_{R}$ representing the feature dimension and $(H,W)$ denoting the BEV resolution.
III-C Coarse Voxel Queries Generator
Figure 3: Illustration of the proposed Coarse Voxel Queries Generator. It combines radar BEV features and multi-view image features to initialize voxel queries with both geometric and semantic priors.
We formulate the 3D voxel queries in space as $\mathcal{Q}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$, where $(H_{V},W_{V},Z_{V})$ denotes the voxel grid resolution. To reduce computational overhead, we set the BEV plane resolution of the voxel grid as $(H_{V},W_{V})=(\frac{H}{2},\frac{W}{2})$. While existing approaches [32, 35] conventionally utilize random initialization for voxel query generation, this methodology potentially introduces additional complexity to the model training process. To address this limitation and enhance the fidelity of view transformation, we introduce a novel initialization method that integrates geometric priors derived from 4D radar data with semantic features extracted from images. This integration enables the generation of coarse-grained voxel queries with both geometric and semantic priors, establishing a more robust foundation for subsequent refinement procedures. Inspired by [25, 34], we design a voxel query initialization pipeline as followings.
In the radar feature processing phase, we initially transform the radar BEV features $\mathcal{F}_{R}$ through Bilinear Interpolation to align with the voxel grid, yielding $\mathcal{F}_{R}^{{}^{\prime}}\in\mathbb{R}^{C_{R}\times H_{V}\times W_{V}}$. Subsequently, we further optimize the feature channels using a Conv-BN-ReLU (CBR). By applying a simple ”unsqueeze” operation to expand the 2D BEV features along the height dimension, we obtain the radar 3D voxel features $\mathcal{Q}_{R}$, which can be mathematically expressed as:
$\mathcal{Q}_{R}=\mathtt{Unsqueeze}(\mathtt{CBR}(\mathcal{F}_{R}^{{}^{\prime}}),Z_{V})$ (1)
For image feature processing, we adopt a methodology similar to [25, 78]. We first define 3D reference points $\mathcal{P}_{ref}\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times 3}$ within the ego-vehicle coordinate frame, based on the shape of the 3D voxel queries. Concurrently, we initialize the voxel features $\mathcal{Q}_{I}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$ to zero. The transformation matrix $\mathbf{T}_{e\to I}\in\mathbb{R}^{3\times 4}$ from the ego-vehicle coordinate frame to the image pixel coordinate is then computed using the camera’s intrinsic matrix $K\in\mathbb{R}^{3\times 4}$ and extrinsic matrix $\mathbf{T}_{e\to c}\in\mathbb{R}^{4\times 4}$:
$\mathbf{T}_{e\to I}=K\mathbf{T}_{e\to c}$ (2)
Leveraging $\mathbf{T}_{e\to I}$, we project reference points onto the each image planes to obtain their corresponding coordinates $(x,y,z)$ on the feature maps. Valid points are determined by two criteria: $(x,y)$ must lie within the feature map boundaries and $z$ must be positive. The feature sampling process employs nearest neighbor interpolation, with a ”last-update” strategy resolving overlapping multi-view regions. The final coarse-grained voxel queries are obtained through element-wise addition:
$\mathcal{Q}=\mathcal{Q}_{R}+\mathcal{Q}_{I}$ (3)
III-D Voxel Queries Encoder
To enhance and refine voxel queries, we employ an $L$-layer Transformer-based architecture for feature encoding. Inspired by [26, 21, 32], we adopt deformable attention for cross-view feature aggregation, which not only alleviates occlusion and ambiguity issues but also improves efficiency by reducing training time.
In the cross-view attention module, the inputs consist of voxel queries $\mathcal{Q}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$, corresponding 3D reference points $\mathcal{P}_{ref}\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times 3}$, and image features $\mathcal{F}_{I}=\{F_{I}^{i}\}_{i=1}^{N_{C}}$. The 3D reference points are projected into 2D views using camera parameters, and image features are sampled and weighted from the hit views. The output features $\mathbf{Q}_{O}$ can be expressed as:
$\mathbf{Q}_{O}^{P}=\frac{1}{|\mathcal{V}_{\text{hit}}|}\sum_{i\in\mathcal{V}_{\text{hit}}}\mathtt{DeformAttn}(\mathbf{Q}^{p},\mathtt{Proj}(\mathcal{P}_{ref}^{p}\ ,\ i),F_{I}^{i})$ (4)
where $i$ denotes the image view index, $\mathbf{Q}^{p}$ and $\mathbf{Q}_{O}^{P}$ represent the $p$-th voxel feature and its output feature respectively, $\mathtt{Proj}(\mathcal{P}_{ref}^{p},i)$ denotes the projection function that maps 3D reference points to the $i$-th image view, and $\mathcal{V}_{\text{hit}}$ represents the set of visible views. Similar to [32], we employ 3D convolutions for neighboring voxel feature interaction instead of computationally expensive 3D self-attention.
III-E Dual-branch Temporal Encoder
Temporal information plays a crucial role in perception systems. Existing methods [21, 78] have demonstrated that leveraging temporal features can effectively address occlusion issues, enhance scene understanding, and improve the accuracy of motion state estimation. However, these approaches are limited to temporal modeling in a single feature space, making it challenging to to capture comprehensive spatio-temporal representations.
To address this limitation, we propose a novel Dual-branch Temporal Encoder module that processes multi-modal temporal features in parallel across BEV and voxel spaces, as shown in Fig. 4. Specifically, the radar BEV branch excels at capturing global geometric features while the image voxel branch focuses on preserving fine-grained semantic information. This complementary dual-branch design not only provides diverse representational capabilities in feature expression and temporal modeling but also achieves an optimized balance between computational cost and feature expressiveness. Additionally, the feature redundancy mechanism significantly enhances the robustness of the perception system.
Figure 4: Illustration of the proposed Dual-branch Temporal Encoder. To eliminate misalignment caused by ego motion, ego poses are used to warp the 2D and 3D reference points and sample features to the current frame. Historical features are then merged using ResNet2D/3D blocks to reduce cross-frame feature interaction and enhance efficiency. To mitigate the impact of moving objects, 2D and 3D deformable attention mechanisms are employed to adaptively fuse features from the current and historical frames.
In temporal feature fusion, a key challenge is feature misalignment caused by ego-motion and dynamic object movement. To address the feature displacement induced by ego-motion, we propose a pose transformation-based feature alignment strategy that precisely aligns historical features with the current frame. Specifically, in the voxel temporal branch, given 3D reference points $\mathcal{P}_{t}^{3D}\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times 3}$ at current frame $t$ in the ego-vehicle coordinate system, we utilize pose sequences from both current and historical frames $\{\mathbf{T}_{t},\mathbf{T}_{t-1},\mathbf{T}_{t-2},...,\mathbf{T}_{t-k}|\mathbf{T}_{t-i}\in\mathbb{R}^{4\times 4}\}$ to warp these reference points to corresponding historical timestamps. Subsequently, we employ trilinear interpolation to sample the temporally aligned features. The detailed computation process is as follows:
$\mathcal{P}_{t-i}^{3D}=\mathbf{T}_{t-i}^{-1}\cdot\mathbf{T}_{t}\cdot\mathcal{P}_{t}^{3D}$ (5)
$\mathbf{F}_{(t-i)\rightarrow t}^{vox}=\mathtt{GridSample3D}(\mathbf{F}_{t-i}^{vox},\mathcal{P}_{t-i}^{3D})$ (6)
where $\mathcal{P}_{t-i}^{3D}$ represents the transformed 3D reference points at historical timestamp $t-i$, and $\mathbf{F}_{(t-i)\rightarrow t}^{vox}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$ denotes the aligned historical features, forming a temporal feature sequence $\{\mathbf{F}_{(t-1)\rightarrow t}^{vox},\mathbf{F}_{(t-2)\rightarrow t}^{vox},...,\mathbf{F}_{(t-k)\rightarrow t}^{vox}\}$. Similarly, for the BEV temporal branch, we define 2D reference points $\mathcal{P}_{t}^{2D}\in\mathbb{R}^{H\times W\times 3}$ by setting the height dimension to zero. The historical features are obtained through bilinear interpolation after warping these reference points to corresponding timestamps, which can be formulated as:
$\mathcal{P}_{t-i}^{2D}=\mathbf{T}_{t-i}^{-1}\cdot\mathbf{T}_{t}\cdot\mathcal{P}_{t}^{2D}$ (7)
$\mathbf{F}_{(t-i)\rightarrow t}^{BEV}=\mathtt{GridSample2D}(\mathbf{F}_{t-i}^{BEV},\mathcal{P}_{t-i}^{2D})$ (8)
where $\mathcal{P}_{t-i}^{2D}$ represents the transformed 2D reference points at historical timestamp $t-i$, and $\mathbf{F}_{(t-i)\rightarrow t}^{BEV}$ denotes the aligned historical features, forming a temporal feature sequence ${\mathbf{F}_{(t-1)\rightarrow t}^{BEV},\mathbf{F}_{(t-2)\rightarrow t}^{BEV},...,\mathbf{F}_{(t-k)\rightarrow t}^{BEV}}$ for the BEV branch.
To further mitigate feature misalignment caused by dynamic objects, we employ deformable attention to adaptively fuse features between current and historical frames. For the voxel temporal branch, we first concatenate the aligned historical features and process them through a simple Res3D block for efficient feature integration, which can be formulated as:
$\mathbf{F}_{hist}^{vox}=\mathtt{ResBlock3D}(\mathtt{Concat}({\mathbf{F}_{(t-1)\rightarrow t}^{vox},\ ...\ ,\mathbf{F}_{(t-k)\rightarrow t}^{vox}}))$ (9)
Subsequently, we employ deformable attention to adaptively integrate current and historical features. The fusion process can be formulated as:
$\mathbf{F}_{O}^{vox_{p}}=\sum_{V\in\{\mathbf{F}_{t}^{vox},\mathbf{F}_{hist}^{vox}\}}\mathtt{DeformAttn3D}(\mathbf{F}^{vox_{p}},p^{3D}\ ,V)$ (10)
where $\mathbf{F}^{vox_{p}}$ and $\mathbf{F}_{O}^{vox_{p}}$ denotes the voxel feature and its output feature located at $p^{3D}=(x,y,z)$, respectively.
For the BEV temporal branch, a similar process is applied. The historical BEV features are first concatenated and processed through a Res2D block:
$\mathbf{F}_{hist}^{BEV}=\mathtt{ResBlock2D}(\mathtt{Concat}({\mathbf{F}_{(t-1)\rightarrow t}^{BEV},\ ...\ ,\mathbf{F}_{(t-k)\rightarrow t}^{BEV}}))$ (11)
Then, deformable attention is employed for feature fusion:
$\mathbf{F}_{O}^{BEV_{p}}=\sum_{V\in\{\mathbf{F}_{t}^{BEV},\mathbf{F}_{hist}^{BEV}\}}\mathtt{DeformAttn2D}(\mathbf{F}^{BEV_{p}},p^{2D},V)$ (12)
where $\mathbf{F}^{BEV_{p}}$ and $\mathbf{F}_{O}^{BEV_{p}}$ denotes the BEV feature and its output feature located at $p^{2D}=(x,y)$, respectively.
This comprehensive approach ensures robust temporal feature fusion by combining global geometric patterns from the BEV branch and fine-grained semantic details from the voxel branch, leading to more accurate perception results.
III-F Cross-Modal BEV-Voxel Fusion Module
Figure 5: Illustration of the proposed Cross-Modal BEV-Voxel Fusion module. Complementary information from both modalities is adaptively fused to generate BEV and voxel features for downstream task decoding. Additionally, two auxiliary tasks which predict the occupied/non-occupied binary 3D occupancy probability and foreground/background BEV segmentation mask, are incorporated to enhance the quality of the generated 3D occupancy and BEV features.
To effectively leverage the temporally-enhanced features from both voxel and BEV spaces, we propose a Cross-Modal BEV-Voxel Fusion module that generates geometrically and semantically rich multi-modal representations for downstream multi-task decoding. As illustrated in Fig. 5, this module adaptively fuses heterogeneous features through attention-weighted mechanisms while employing auxiliary tasks to further enhance the quality of generated features.
Specifically, the module first upsamples the low-resolution voxel features through a 3D deconvolution block to obtain high-resolution features $\mathbf{F}^{vox}\in\mathbb{R}^{C\times H\times W\times Z}$ for subsequent fusion. For voxel feature enhancement, the radar BEV features $\mathbf{F}^{BEV}$ are first processed through Conv-BN-ReLU blocks in 2D to reshape the feature channels, followed by an unsqueeze operation that expands the 2D BEV features along the height dimension. The expanded features are then concatenated with voxel features and processed through convolutional blocks to reduce channel dimensions. Finally, a residual structure with attention mechanism is employed to obtain the fused features. This process can be formulated as:
$\mathbf{F}_{temp}^{vox}=f_{3D}(\mathtt{Concat}[\mathbf{F}^{vox},\mathtt{Unsqueeze}(f_{2D}(\mathbf{F}^{BEV}))])$ (13)
$\mathbf{F}_{fused}^{vox}=\mathbf{F}^{vox}+\sigma(g_{3D}(\mathbf{F}_{temp}^{vox}))\odot\mathbf{F}_{temp}^{vox}$ (14)
where $f_{2D}(\cdot)$ and $f_{3D}(\cdot)$ represent the Conv-BN-ReLU blocks in 2D and 3D, $g_{3D}(\cdot)$ represents the convolutional operations in the attention block, $\sigma(x)=1/(1+e^{-x})$ is the sigmoid function, and $\odot$ denotes element-wise multiplication. $\mathbf{F}_{temp}^{vox}$ and $\mathbf{F}_{fused}^{vox}$ are intermediate and final fused voxel features.
Similarly, for BEV feature enhancement, the voxel features $\mathbf{F}^{vox}$ are first compressed along the height dimension through a squeeze operation, followed by CBR2D blocks to adjust feature channels, obtaining transformed features $\mathbf{F}^{BEV^{\prime}}\in\mathbb{R}^{C\times H\times W}$. The processed features are then concatenated with radar BEV features and refined through convolutional blocks. A similar residual attention structure is employed to obtain the final fused BEV features. This process can be formulated as:
$\mathbf{F}_{temp}^{BEV}=f_{2D}(\mathtt{Concat}[\mathbf{F}^{BEV},f_{2D}(\mathtt{Squeeze}(\mathbf{F}^{vox}))])$ (15)
$\mathbf{F}_{fused}^{BEV}=\mathbf{F}^{BEV^{\prime}}+\sigma(g_{2D}(\mathbf{F}_{temp}^{BEV}))\odot\mathbf{F}_{temp}^{BEV}$ (16)
where $g_{2D}(\cdot)$ represents the convolutional operations in the 2D attention block, and $\mathbf{F}_{temp}^{BEV}$ and $\mathbf{F}_{fused}^{BEV}$ are intermediate and final fused BEV features, respectively.
Additionally, to enhance the final feature representations and improve feature quality for subsequent decoding, we incorporate auxiliary tasks with explicit supervision. For the fused voxel features, we estimate binary occupancy masks $\mathbf{M}^{vox}\in\mathbb{R}^{H\times W\times Z}$ through 3D convolutional blocks, supervised by binary occupied/non-occupied ground truth derived from semantic occupancy labels. Similarly, for the fused BEV features, we project detection ground truth onto the BEV plane to obtain binary segmentation ground truth representing foreground objects and background, and generate corresponding prediction mask $\mathbf{M}^{BEV}\in\mathbb{R}^{H\times W}$ through 2D convolutional blocks. Both auxiliary tasks are supervised using a combination of Dice and Binary Cross-Entropy (BCE) losses. These auxiliary binary supervision signals help guiding the feature learning process, ensure the generation of more discriminative features progressively in the suitable area for downstream tasks with more specific semantic information, i.e., 3D semantic occupancy prediction and 3D object detection.
III-G Multi-task Training
Based on the enhanced geometric and semantic-aware BEV and voxel representations, we can perform end-to-end training for joint 3D object detection and occupancy prediction, enabling comprehensive perception capabilities.
For 3D object detection, following [21], we use DETR-based head unless specified otherwise. The head takes the fused BEV features $\mathbf{F}_{fused}^{BEV}$ as input and directly predicts object categories and attributes. Specifically, each 3D bounding box is parameterized by 10 parameters: dimensions $(l,w,h)$, center location $(x_{o},y_{o},z_{o})$, orientation $(\cos{\theta},\sin{\theta})$, and velocity $(v_{x},v_{y})$. This end-to-end approach eliminates the need for post-processing steps such as NMS. The detection loss $\mathcal{L}_{det}$ consists of a focal loss for classification and an L1 loss for regression, which can be formulated as:
$\mathcal{L}_{det}=\lambda_{1}\mathcal{L}_{cls}+\lambda_{2}\mathcal{L}_{reg}$ (17)
where $\lambda_{1}=2.0$ and $\lambda_{2}=0.25$ are hyperparameters to balance the classification and regression losses.
For occupancy prediction, we employ a simple MLP on the fused voxel features $\mathbf{F}_{fused}^{vox}$ to predict semantic occupancy for each voxel. The occupancy prediction loss $\mathcal{L}_{occ}$ consists of three components: a primary cross-entropy loss $\mathcal{L}_{ce}$ for basic supervision, and two affinity losses $\mathcal{L}_{scal}^{geo}$ and $\mathcal{L}_{scal}^{sem}$ proposed by [33] to optimize scene-wise and class-wise metrics, respectively. The occupancy loss $\mathcal{L}_{occ}$ can be formulated as:
$\mathcal{L}_{occ}=\mathcal{L}_{ce}+\mathcal{L}_{scal}^{geo}+\mathcal{L}_{scal}^{sem}$ (18)
For both foreground/background BEV segmentation and occupied/non-occupied binary occupied prediction auxiliary tasks, we employ a combination of BCE loss and Dice loss for supervision. The auxiliary loss $\mathcal{L}_{aux}$ can be formulated as:
$\mathcal{L}_{aux}=\mathcal{L}_{bce}^{occupied}+\mathcal{L}_{dice}^{occupied}+\mathcal{L}_{bce}^{seg}+\mathcal{L}_{dice}^{seg}$ (19)
The total loss is formulated as:
$\mathcal{L}_{total}=\mathcal{L}_{det}+\mathcal{L}_{occ}+\mathcal{L}_{aux}$ (20)
IV Experiments and Performance Analysis
IV-A Dataset and Evaluation Metrics
IV-A1 Dataset
We evaluate our method on three datasets: OmniHD-Scenes [13], VoD [14], and TJ4DRadSet [15]. All datasets provide data measured by synchronized sensors including 4D imaging radar, camera, and LiDAR, along with 3D object annotations. Notably, OmniHD-Scenes also provides semantic labels for static scenes and dense 3D occupancy ground truth, enabling comprehensive evaluation of our multi-task framework.
The OmniHD-Scenes dataset is a recent large-scale multi-modal dataset featuring a comprehensive omnidirectional sensor suite, which consists of six cameras, six 4D imaging radars, and a 128-beam LiDAR. The dataset contains 1501 sequences with 30-seconds for each, captures under diverse scenarios including nighttime, rainy conditions, and complex traffic situations. Currently, annotations are provided for 200 clips, comprising 3D tracking annotations, static scene segmentation annotations, and dense semantic occupancy ground truth. The 3D bounding box annotations cover four object categories: cars, pedestrians, riders, and large vehicles, while the occupancy ground truth includes 11 semantic classes. The annotated data contains 11921 keyframes in total, with 8321 frames for training and 3600 frames for testing.
The VoD dataset was collected from the campus, suburbs, and old town in Delft, with 5139 training frames and 1296 validation frames. The TJ4DRadSet dataset has 5717 frames for training and 2040 frames for testing, including diverse road scenarios such as urban streets, elevated highways, and industrial areas. Both of them provide synchronized sensor data from a forward-facing setup, consisting of a 4D imaging radar, camera, and LiDAR, along with 3D bounding box annotations. The former contains three categories: cars, pedestrians, and cyclists, while the latter also supplies trucks as supplement.
IV-A2 Evaluation Metrics
For the OmniHD-Scenes dataset, we utilize officially defined metrics to evaluate the performance of 3D detection and occupancy prediction within a detection area of ±60m longitudinally and ±40m laterally around the ego vehicle. For 3D detection, we employ mean Average Precision (mAP) along with four mean True Positive metrics (mTP): mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Additionally, we adopt the OmniHD-Scenes Detection Score (ODS) to assess comprehensive performance, as:
$ODS=\frac{1}{8}[4mAP+\sum_{mTP\in TP}(1-\min(1,mTP))]$ (21)
For 3D occupancy prediction evaluation, we employ two key metrics: mean Intersection over Union (mIoU) for semantic accuracy and scene completion IoU for geometric accuracy. The mIoU is calculated by averaging IoU scores across all semantic categories, where IoU measures the overlap between predicted and ground truth occupancy states for each category. The scene completion IoU specifically assesses the model’s capability to distinguish between free and occupied spaces by computing IoU metrics for both spatial states. The details are expressed as follows:
$\mathrm{mIoU}=\frac{1}{C}\sum_{i=1}^{C}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ (22)
$\mathrm{IoU}=\frac{TP}{TP+FP+FN}$ (23)
For both VoD and TJ4DRadSet datasets, we evaluate 3D detection performance using Average Precision (AP) and mean Average Precision (mAP) by following the official evaluation protocols [14, 15].
IV-B Implementation Details
For the OmniHD-Scenes dataset, we constrain the point cloud range to $(-60,60)$ m, $(-40,40)$ m, and $(-3,5)$ m along the X-, Y-, and Z-axes, respectively. The radar input consists of 3-frame accumulated point clouds, with each point characterized by a feature vector $[x,y,z,Power,SNR,v_{xr},v_{yr}]$, where $Power$ and $SNR$ represent amplitude and signal-to-noise ratio, while $v_{xr}$ and $v_{yr}$ denote the compensated absolute velocity components. All six camera images are resized to $544\times 960$. The low-resolution voxel query dimensions $H_{V}\times W_{V}\times Z_{V}$ are set to $80\times 120\times 8$, the BEV feature map size $H\times W$ is $160\times 240$, and the occupancy ground truth resolution is $160\times 240\times 16$. We employ a DETR-based detection head with 900 object queries and retain the top 300 predicted boxes with highest confidence scores during inference. For both VoD and TJ4DRadSet datasets, we followed exact same settings as most of the state-of-the-arts [68, 70, 71].
Our model implementation is based on the MMDetection3D [79] framework and trained using NVIDIA GeForce RTX 4090D GPUs. For the camera encoder, we utilize pre-trained weights from FCOS3D [16] for backbone, maintaining consistency with [21]. The 4D radar encoder, inherited from RadarPillarNet [68], is trained from scratch for 3D detection on their respective datasets. We employ the AdamW optimizer for training, with different learning rate for each dataset: for OmniHD-Scenes, we train for 16 epochs with a learning rate of $2\times 10^{-4}$; for VoD and TJ4DRadSet, we train for 16 and 20 epochs respectively, both with a learning rate of $1\times 10^{-4}$.
All ablation experiments are conducted on the OmniHD-Scenes dataset under a multi-task learning setting with two temporal frames, unless otherwise specified.
IV-C Performance Comparison with State-of-the-arts
Figure 6: Qualitative results of Doracamom on the OmniHD-Scenes test set. The left column shows predicted 3D bounding boxes projected onto six camera views, where cars, pedestrians, riders, and large vehicles are displayed in yellow, blue, green, and pink, respectively. The middle column presents the BEV perspective, with black points representing LiDAR data and red points indicating 4D radar data, where ground truth and predicted boxes are shown in green and blue. The right column shows the predicted occupancy grids and ground truth occupancy grids, using the same color scheme as Table II. These results demonstrate that Doracamom achieves robust multi-task perception performance across diverse scenarios. Please zoom in to see more detail.
Figure 7: Visualization of BEV feature maps from different methods. Compared with BEVFusion [80] and BEVFormer [21], Doracamom generates feature maps with clearer object contours and more prominent features.
Figure 8: Qualitative results of Doracamom on the TJ4DRadSet and VoD datasets. For each sample, the upper row shows predicted 3D bounding boxes projected onto camera views, and the lower row presents BEV visualization with red points representing 4D radar data and blue boxes indicating detection results.
IV-C1 Results on OmniHD-Scenes
TABLE I: Comparison of state-of-the-art approach with ours, for 3D object detection task on the OmniHD-Scenes test set. “L” denotes LiDAR, “C” denotes camera and “R” denotes 4D radar. ”-128” and ”-32” represent 128 lines and manually downsampled 32 lines, respectively. The last four columns show the AP for each object type. “Ped.” and “LVeh.” represent pedestrian and large vehicle, respectively. Bold and underline denote the first and the second best performances among all the approaches without using LiDAR. Doracamom-S does not utilize temporal information.
[…]
PointPillars-128 [3] (CVPR 2019) - L - 61.15 55.54 0.2825 0.1980 0.5223 1.8763 85.43 28.63 69.18 61.36
PointPillars-32 [3] (CVPR 2019) - L - 57.24 52.66 0.3040 0.2038 0.5692 1.8731 82.74 24.22 65.86 56.16
PointPillars [3] (CVPR 2019) - R - 23.82 37.21 0.6752 0.2447 0.3776 0.6789 52.74 0.69 28.57 13.29
RadarPillarNet [68] (IEEE T-IM 2023) - R - 24.88 37.81 0.6597 0.2389 0.3736 0.6982 52.99 2.06 29.45 15.02
LSS-Depth [20] (AAAI 2023) 544×960 C R50 22.44 26.01 1.0238 0.2230 0.5942 2.0138 50.42 4.37 24.56 10.42
BEVformer [21] (ECCV 2022) 544×960 C R50 26.49 28.10 1.1430 0.2315 0.5799 1.6666 50.74 11.69 30.42 13.11
BEVformer-T [21] (ECCV 2022) 544×960 C R50 29.17 30.54 1.1046 0.2346 0.4889 1.0797 53.64 14.48 33.55 15.01
BEVformer [21] (ECCV 2022) 864×1536 C R101-DCN 30.10 30.55 1.0633 0.2266 0.5331 1.6625 55.40 14.05 34.24 16.70
BEVformer-T [21] (ECCV 2022) 864×1536 C R101-DCN 32.22 32.57 1.0637 0.2271 0.4558 1.0683 57.61 17.37 38.02 15.87
PanoOcc [35] (CVPR 2024) 544×960 C R50 29.17 28.55 1.1500 0.2446 0.6378 1.6066 51.58 15.82 35.02 14.26
BEVFusion [80] (NeurIPS 2022) 544×960 C&R R50 33.95 43.00 0.5730 0.2165 0.3814 0.7474 56.25 11.66 50.90 16.99
RCFusion [68] (IEEE T-IM 2023) 544×960 C&R R50 34.88 41.53 0.5676 0.2135 0.3711 0.9208 57.17 12.87 51.35 18.11
Doracamom-S (ours) 544×960 C&R R50 37.60 41.31 0.6724 0.2329 0.4359 0.8579 58.94 17.84 52.72 20.89
Doracamom (ours) 544×960 C&R R50 39.12 46.22 0.6646 0.2331 0.3545 0.6151 61.12 19.83 53.35 22.18
3D Object Detection. Table I presents the performance comparison of different methods on the OmniHD-Scenes test set for 3D detection tasks. Our proposed Doracamom achieves superior overall performance (39.12 mAP & 46.22 ODS) compared to other approaches based on 4D radar, camera, or their fusion. Specifically, it outperforms BEVFusion [80] by +5.17 mAP and +3.22 ODS, while surpassing RCFusion [68] by +4.24 mAP and +4.69 ODS. Even in the single-frame setting without the DTE module, our model outperforms all other methods in terms of mAP. Furthermore, Doracamom significantly narrows the performance gap with LiDAR-based PointPillars [3] (46.22 ODS vs. 55.54 ODS), which demonstrates both the effectiveness of our proposed architecture and the tremendous potential of low-cost sensor configurations in autonomous driving perception systems. In terms of TP metrics, our method achieves best performance in both mAOE and mAVE, reaching 0.3545 and 0.6151, respectively. Notably, benefiting from the Doppler information inherent in 4D radar, approaches leveraging 4D radar independently or incorporating 4D radar-camera fusion demonstrate remarkable velocity estimation accuracy compared to camera-only or LiDAR-based methods, exhibiting substantially lower errors in velocity predictions. Furthermore, Doracamom achieves even more precise velocity estimation through efficient exploitation of 4D radar features and the DTE module. For each object category, our method achieves the highest AP scores across all classes: Car, Pedestrian, Rider, and Large Vehicle. Notably, all methods exhibit relatively low detection accuracy for pedestrians and large vehicles, which can be attributed to several challenging factors in the dataset. The dataset contains numerous crowded scenes with dozens of people in small areas, leading to severe occlusions. Moreover, the features extracted from both image and radar data are often incomplete and less distinctive. Additionally, pedestrians occupy only a few grids in the BEV features, further increasing the detection difficulty. For Large Vehicles, their substantial size often results in incomplete contours in both radar and camera views, leading to significant size discrepancies.
The visualization results shown in Fig. 6 denote that Doracamom can provide reliable performance across both day and night scenarios. It achieves high detection accuracy in crowded and complex scenes, with only occasional missed detections of distant occluded objects. Fig. 7 illustrates the BEV feature maps of different methods. It can be observed that the feature map of Doracamom displays distinct object boundaries and highly distinguishable features, with no significant issues such as severe stretching or distortion of the objects.
TABLE II: Comparison of state-of-the-art approach with ours, for 3D occupancy prediction task on the OmniHD-Scenes test set. ”C” denotes camera and ”R” denotes 4D radar. The last eleven columns show the IoU for each semantic type. Bold and underline denote the first and the second best performances. Doracamom-S does not utilize temporal information.
Methods Image Res. Modality Backbone SC IoU mIoU
[…]
C-CONet [81] (ICCV 2023) 544×960 C R50 25.69 13.42 20.03 3.51 11.71 16.62 0.79 1.14 22.75 33.57 14.82 17.73 4.93
SurroundOcc [32] (ICCV 2023) 544×960 C R50 28.61 15.20 21.46 3.96 10.76 16.58 1.57 2.99 21.63 48.52 18.31 16.73 4.71
BEVFormer [21] (ECCV 2022) 544×960 C R50 27.04 14.97 20.64 5.87 14.40 16.68 1.52 3.64 20.64 46.61 16.19 14.80 3.69
BEVFormer-T [21] (ECCV 2022) 544×960 C R50 28.42 16.23 22.73 5.45 14.70 18.21 3.09 3.87 21.54 48.15 17.58 17.77 5.48
PanoOcc [35] (CVPR 2024) 544×960 C R50 26.36 15.20 22.42 5.91 13.58 17.98 3.11 3.36 21.46 50.47 15.90 11.20 1.80
BEVFormer [21] (ECCV 2022) 864×1536 C R101-DCN 28.30 16.41 23.72 6.37 16.33 20.44 1.78 3.78 22.21 48.55 17.88 15.49 3.99
BEVFormer-T [21] (ECCV 2022) 864×1536 C R101-DCN 29.74 17.49 24.90 6.48 16.45 21.49 2.87 4.62 22.51 49.92 18.59 18.53 5.96
BEVFusion [80] (NeurIPS 2022) 544×960 C&R R50 27.02 16.24 27.02 4.78 21.71 21.59 1.55 2.78 25.21 44.35 12.32 13.06 4.25
M-CONet [81] (ICCV 2023) 544×960 C&R R50 27.74 16.08 25.21 3.42 17.53 21.46 0.88 0.58 29.88 34.48 14.89 19.57 8.98
Doracamom-S (ours) 544×960 C&R R50 31.46 19.49 30.10 6.71 23.60 24.31 2.85 6.55 25.77 49.72 16.53 19.57 8.72
Doracamom (ours) 544×960 C&R R50 33.96 21.81 30.81 7.22 24.33 24.70 4.49 7.84 34.49 52.00 20.86 21.68 11.49
TABLE III: Comparison of state-of-the-art approach with ours, for 3D object detection task under adverse conditions (night and rainy weather) on the OmniHD-Scenes test set.
[…]
BEVformer-R101 [21] C 28.11 29.36 26.66 15.03
BEVformer-T-R101 [21] C 30.39 31.61 28.41 15.84
PanoOcc [35] C 26.09 27.16 24.02 14.02
BEVFusion [80] C&R 35.83 44.95 25.32 15.36
M-CONet [81] C&R 26.73 15.30
Doracamom-S (ours) C&R 38.75 43.47 29.94 18.81
Doracamom (ours) C&R 41.86 48.74 31.06 20.30
3D Semantic Occupancy. Table II presents the performance comparison of different methods on the OmniHD-Scenes validation set for occupancy prediction tasks. Our proposed Doracamom achieves superior overall performance (33.96 SC IoU & 21.81 mIoU) compared to other approaches. When BEVFormer [21] utilizes a larger backbone network (R101-DCN) and higher resolution image input (864*1536), its performance surpasses multi-sensor fusion methods like M-CONet [81] that combines camera and 4D radar data. Nevertheless, with our well-designed architecture, even without using temporal information, Doracamom-S significantly outperforms BEVFormer-T by +1.72 SC IoU and +2.00 mIoU. Moreover, examining the IoU metrics for each category, we observe that both Doracamom and Doracamom-S achieve superior performance compared to other models in detecting foreground objects of particular interest for object detection, specifically cars, pedestrians, riders, and large vehicles. This performance advantage demonstrates that in a multi-task setting, object detection and occupancy prediction can work synergistically to facilitate both tasks. This significant advantage is also visually apparent in Fig. 6, showing that Doracamom maintains strong occupancy prediction performance in both day and night scenarios. In some cases, it even successfully completes ground holes present in the ground truth. However, the model still exhibits some prediction errors in occluded regions.
Performance under Adverse Conditions. Table III denotes the performance of different models under adverse conditions, where Doracamom achieves better results with 41.86 mAP and 48.74 ODS, consistently outperforming other methods and showing stronger robustness. Compared to the results in Table I, camera-based methods show decreased detection accuracy, while methods combining 4D radar and camera demonstrate improved performance. Doracamom also achieves 31.06 SC IoU and 20.30 mIoU in occupancy prediction tasks, consistently outperforming other models, which indicates its superior robustness. However, compared to Table II, all models show some performance degradation, highlighting the crucial role of camera input in occupancy prediction, as camera performance deterioration directly impacts overall model performance. Nevertheless, with the assistance of 4D radar, the performance degradation remains within an acceptable range. This indicates that 4D radar provides better environmental adaptability in challenging conditions.
TABLE IV: Comparison of resources consumption and efficiency with other state-of-the-arts (tested on the 4090D GPU).
Method Resource & Efficiency Metrics
[…]
PanoOcc [35] 5.03G 51.94M 5.5 29.17 28.55 26.36 15.20
BEVFusion-OD [80] 7.96G 57.26M 3.6 33.95 43.00
BEVFusion-OCC [80] 7.98G 57.22M 3.2 27.02 16.24
Doracamom-S 4.71G 49.63M 4.8 37.60 41.31 31.46 19.49
Doracamom-2frames 4.72G 52.67M 4.4 38.21 44.52 33.16 21.44
Resources Consumption and Efficiency. Table IV presents a comparison of different models in terms of resource consumption and efficiency. Compared to existing methods, our Doracamom series models demonstrate an excellent balance between performance and efficiency. In terms of resource consumption, Doracamom-S requires only 4.71G memory and 49.63M parameters, making it more lightweight than BEVFusion ($\sim$8G memory and 57M parameters) and PanoOcc (5.03G memory and 51.94M parameters). Even with 2 frames incorporated, Doracamom-2frames maintains relatively low resource usage (4.72G memory, 52.67M parameters). Regarding inference efficiency, Doracamom-S and Doracamom-2frames achieve 4.8FPS and 4.4FPS respectively, significantly outperforming the BEVFusion series (3.2-3.6FPS). While slightly slower than PanoOcc (5.5FPS), our models demonstrate substantial performance advantages: Doracamom-2frames achieves optimal performance across all evaluation metrics, significantly surpassing other methods.
IV-C2 Results on VoD & TJ4DRadSet
TABLE V: Comparison of state-of-the-art approach with ours, for 3D object detection task on VoD validation set and TJ4DRadSet test set, respectively. “C” denotes camera and “R” denotes 4D radar. indicates the use of extra LiDAR data during training procedure. Bold and underline denote the first and the second best performances among all the approaches. To ensure fair comparison, Doracamom-S does not utilize temporal information, same as others.
Method Modality AP3D (%) of VoD AP3D (%) of TJ4DRadSet
Car Ped. Cyclist mAP Car Ped. Cyclist Truck mAP
CenterPoint [4] (CVPR 2021) R 32.74 38.00 65.51 45.42 22.03 25.02 53.32 15.92 29.07
SMURF [65] (IEEE T-IV 2024) R 42.31 39.09 71.50 50.97 28.47 26.22 54.61 22.64 32.99
${}^{\text{\textdagger}}$ [66] (AAAI 2025) R 41.89 43.51 70.83 52.08 - - - - -
RCFusion [68] (IEEE T-IM 2023) C&R 41.70 38.95 68.31 49.65 29.72 27.17 54.93 23.56 33.85
RCBEVDet [40] (CVPR 2024) C&R 40.63 38.86 70.48 49.99 - - - - -
LXL [70] (IEEE T-IV 2024) C&R 42.33 49.48 77.12 56.31 - - - - 36.32
UniBEVFusion [69] (IEEE ICRA 2025) C&R 42.22 47.11 72.94 54.09 44.26 27.92 51.11 27.75 37.76
HGSFusion [72] (AAAI 2025) C&R 51.67 52.64 72.58 58.96 - - - - 37.21
${}^{\text{\textdagger}}$ [71] (IEEE RA-L 2025) C&R 53.16 49.98 76.11 59.75 59.43 26.57 51.30 30.00 41.82
Doracamom-S (ours) C&R 53.35 48.94 76.99 59.76 52.18 33.61 56.38 34.79 44.24
Table V demonstrates the detection performance of various methods on both VoD and TJ4DRadSet datasets. Since existing state-of-the-art methods do not utilize temporal information, we conducted experiments without temporal data to ensure a fair comparison. On the VoD dataset, Doracamom-S achieves superior detection performance with 59.76 mAP, slightly outperforming the current state-of-the-art method SGDet3D [71], despite the latter utilizing additional LiDAR data for depth supervision. Compared to HGSFusion [72] and LXL [70], our method shows improvements of 0.8 mAP and 3.45 mAP respectively. In terms of per-category AP metrics, our method achieves the best performance in car detection and second-best performance in cyclist detection.
On the TJ4DRadSet dataset, Doracamom-S demonstrates an even more substantial lead with an mAP of 44.24, surpassing SGDet3D [71] and HGSFusion [72] by 2.42 mAP and 7.03 mAP respectively. Notably, our method achieves state-of-the-art results across multiple categories including pedestrians, cyclists, and trucks. These results demonstrate that our approach effectively handles both large objects with abundant radar reflection points and smaller objects with fewer reflections, validating the architecture’s superior capability in feature fusion and utilization across complex scenarios.
Fig. 8 presents visualization results from both TJ4DRadSet and VoD datasets. The results show that Doracamom-S maintains excellent performance across diverse scenarios, demonstrating low localization error compared to ground truth and accurate category prediction. Only occasional missed detections are observed under challenging occlusion conditions.
IV-D Ablation Study
TABLE VI: Ablation study on proposed modules.
Method Components Metrics
[…]
Baseline 35.21 38.28 29.42 18.39
+CVQG 37.01 41.48 31.03 19.55
+DTE 37.72 43.87 32.89 21.15
+CMF 38.21 44.52 33.16 21.44
Ablation Study on Main Components. Table VI demonstrates the impact of our core components on model performance, including Coarse Voxel Query Generation (CVQG), Dual-branch Temporal Encoder (DTE), and Cross-modal BEV-Voxel Fusion module (CMF). Starting from a baseline model (random initialization, current-frame features, and feature summation), we progressively add components for analysis.
The baseline model achieves 35.21 mAP and 38.28 ODS in detection tasks, while reaching 29.42 IoU and 18.39 mIoU in occupancy prediction. With the introduction of the VQG module, all metrics show significant improvements (+1.80 mAP, +3.20 ODS, +1.61 IoU, +1.16 mIoU), validating its effectiveness in generating high-quality initial queries. Further incorporating the DTE module (2 frames) enhances detection performance to 37.72 mAP and 43.87 ODS, while occupancy prediction improves to 32.89 IoU and 21.15 mIoU, demonstrating its powerful temporal fusion capabilities. Finally, the complete model with the CMF module achieves optimal performance across all metrics, confirming its strong multi-modal feature fusion ability. These results validate the effectiveness of each proposed component and their synergistic benefits when combined within our framework.
TABLE VII: Ablation study on CVQG.
[…]
(a) 36.92 42.94 32.51 20.30
(b) 37.23 43.27 32.24 20.05
(c) 38.21 44.52 33.16 21.44
Ablation Study on CVQG. Table VII presents the ablation results of the CVQG module, with random initialization serving as the baseline. When incorporating image semantic features for query initialization, detection performance shows slight improvements (+0.31 mAP, +0.33 ODS), but occupancy prediction performance experiences a minor decrease. Furthermore, when integrating both camera semantic features and 4D radar geometric features for initialization, all metrics demonstrate significant improvements. Compared to the baseline, detection performance increases by 1.29 mAP and 1.58 in ODS, while occupancy prediction improves by 0.65 IoU and 1.14 mIoU. These results demonstrate that CVQG, by effectively combining 4D radar geometric and dynamic features with image semantic priors, successfully enhances the performance in both detection and occupancy prediction tasks.
TABLE VIII: Ablation study on proposed DTE modules.
Length Naive ConvBlock Ours
[…]
T = 1 37.60 41.31 31.46 19.49 0.8579 37.60 41.31 31.46 19.49 0.8579
T = 2 38.13 43.96 31.85 20.15 0.6927 38.21 44.52 33.16 21.44 0.6686
T = 3 38.81 45.05 32.57 21.03 0.6712 38.75 45.35 33.55 21.64 0.6576
T = 4 39.56 45.65 32.80 21.51 0.6608 39.12 46.22 33.96 21.81 0.6151
Ablation Study on DTE. Table VIII presents the ablation study results of the DTE module. We use a simple feature concatenation followed by convolution blocks to merge aligned temporal feature sequences as our baseline, termed Naive ConvBlock. The results show that both methods performance improves as the temporal length increases.
When incorporating two-frame temporal information, our proposed DTE module demonstrates stronger performance improvements compared to the naive convolution approach (38.21 mAP vs 38.13 mAP, 44.52 ODS vs 43.96 ODS), with particularly significant gains in occupancy prediction tasks (+1.31 IoU, +1.29 mIoU). With temporal length increased to four frames, although the naive convolution approach achieves optimal mAP, our method obtains the best performance in spatial perception metrics (ODS, IoU, and mIoU). This indicates that the DTE module, through its deformable attention mechanism, can adaptively fuse temporal features, achieving more comprehensive spatio-temporal representations and thus excelling in handling occlusions and enhancing scene understanding. Notably, our method also demonstrates a considerable advantage in velocity estimation (0.6151 vs 0.6608 in mAVE), further validates the effectiveness of the DTE module in spatio-temporal and dynamic object representations.
TABLE IX: Ablation study on proposed CMF modules.
Method Metrics
[…]
Add 37.72 43.87 32.59 21.11
Concat 37.57 44.17 32.94 20.84
Ours w/o Aux. 37.84 44.42 32.81 21.30
Ours w/ Aux. 38.21 44.52 33.16 21.44
Ablation Study on CMF. Table IX presents the ablation study results of the Cross-modal BEV-Voxel Fusion (CMF) module. We compare CMF with two basic feature fusion strategies (feature addition and concatenation). Our proposed CMF module, which adaptively fuses heterogeneous features from BEV and voxel spaces through attention-weighted mechanisms, demonstrates clear advantages even without auxiliary tasks (Ours w/o Aux.): achieving 0.19 improvement in mIoU compared to feature addition, and 0.25 improvement in ODS compared to feature concatenation. With the introduction of auxiliary tasks (Ours w/ Aux.) - foreground/background BEV segmentation and occupied/non-occupied binary prediction - the model achieves comprehensive performance improvements (+0.37 mAP, +0.10 ODS, +0.35 IoU, +0.14 mIoU) over the version without auxiliary tasks. These experimental results not only validate the effectiveness of our adaptive fusion strategy in integrating multi-modal features but also demonstrate that auxiliary tasks can successfully enhance feature representation through additional supervisory signals.
V Conclusion
In this paper, we present Doracamom, the first unified multi-task perception framework with multi-view camera and 4D radar fusion. Specifically, we propose a coarse voxel query generator that initializes geometrically and semantically aware voxel queries to effectively leverage multi-modal features; a dual-branch temporal encoder that adaptively fuses temporal features in both BEV and voxel spaces while considering dynamic and static elements in the environment; and finally, an attention-based cross-modal BEV-voxel fusion module that adaptively integrates heterogeneous features and incorporates auxiliary tasks to address feature ambiguity, generating high-quality feature representations for downstream tasks. Experimental results on three datasets, OmniHD-Scenes, VoD, and TJ4DRadSet, demonstrate that our method achieves state-of-the-art performance in both 3D object detection and 3D semantic occupancy prediction tasks.